OpenStudy (anonymous):

√(2-(-1))^2+(-1-2)^2 I know the answer is 4.24, but I don't know how to get there

OpenStudy (amorfide):

\[\sqrt{(2-(-1))^{2}+(-1-2)^{2}}\] okay so 2--1=2+1=3 -1-2=3 so you get \[\sqrt{(3)^{2}+(-3)^{2}}\]

OpenStudy (amorfide):

you can finish this I am sure

OpenStudy (amorfide):

-1-2=-3 sorry typo of the sign

OpenStudy (anonymous):

But they cancel each other out right?

OpenStudy (anonymous):

9+(-9)= 0 so I just end up with √0

Nnesha (nnesha):

\[\huge\rm (-3)^2 = ??\]

OpenStudy (anonymous):

Thanks I got it

Nnesha (nnesha):

if you have even even exponent of negative number ansawer always going to be POSITIVE and if u have odd number of exponent of negative number answer always going to be NEGATIVE

OpenStudy (amorfide):

sorry openstudy is always so slow I only just got the notification of the reply thank you @Nnesha

Nnesha (nnesha):

np :) btw i didnb't do anything thanks to you :) you helped her not mehh